Q.6.2   The financial manager of a national warehouse chain has found from previous experience that 45% of clients pay by credit card, 35% pay on account, while the remainder pay using cash. He wishes to know if this is still the case. A random sample of 1000 sales receipts from the last month produced the following results: (8).

Payment Method

Credit card

Account

Cash

Total

Number of receipts

500

315

185

1000

Using a 1% significance level, test whether the payment method pattern has changed.
Answer #1
observed frequencey, O expected proportion expected frequency,E (O-E)²/E
500 0.450 450.00 5.556
315 0.350 350.00 3.500
185 0.200 200.00 1.125

Ho: two variables are independent of each other
H1: two variables are not not independent to each other

chi square test statistic,X² = Σ(O-E)²/E =   10.181              
                  
level of significance, α=   0.01              
Degree of freedom=k-1=   3   -   1   =   2
                  
P value =   0.0062   [ excel function: =chisq.dist.rt(test-stat,df) ]          
Decision: P value < α, Reject Ho                  

Hence the payment method pattern has changed.

THANKS

revert back for doubt

please upvote